Shift report 20000809 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 09 2000 - 15:54:00 EDT

  • Next message: Konstantin Olchanski: "Re: Shift report 20000809 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Aug  9 15:53:59 2000, Shift: 08:00-16:00
    Supervisor: HB, with: B1,CE,ER
    Log book#4 pages 41 - 42
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam all day. False firealarms in our hall 
        (MCR will check tomorrow) - then QP Switch problem)
    2. Ramiro changed flow transducer for FFS
        (no HV allowed before 17.00)
    3. Konstantin fixed the DAQ problem (started previous shift
         at 1:45 fixed at 15:00)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 09 2000 - 15:57:17 EDT